OpenStudy (anonymous):

what is the slope between -4,1 and -2,-6

OpenStudy (igreen):

To find the slope between two points, use the slope equation: \(m = \dfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (igreen):

In this case: \(y_2 = -6\) \(y_1 = 1\) \(x_2 = -2\) \(x_1 = -4\) Plug them in: \(m = \dfrac{-6-1}{-2+4}\) \(m = \dfrac{-7}{2}\) So the slope is \(\dfrac{-7}{2}\) or \(-3.5\).

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @markwiggins